If 15c3r = 15cr+3, then r =

1. 2
2. 3
3. 4
4. 5

Answer by farohw(175)


Hello,

I remember this from Discrete Math.

15 is a superscript of C and 3r, r+3 are subscripts of C. If 15C3r = 15Cr+3, then we follow the formula:

nCx = nCy where x = y OR x + y = n.

Let n represent 15, x = 3r and y = r +3:

3r = r + 3 and 3r + r + 3 = 15

3r = r + 3 ---> 3r - r =3 then 2r = 3 which is r = 3/2

3r + r + 3 = 15 ---> 4r = 12 which is r = 3

r cannot be a fraction since this will give us 1.5 therefore the answer will be r = 3.